OpenStudy (anonymous):

4(7x - 1) = 3(2x + 1) - 4(x + 6) i got -11/5 but its not right

OpenStudy (anonymous):

try again

OpenStudy (anonymous):

28x-4=6x+3-4x-24 26x=-17 x=-17/26

OpenStudy (anonymous):

x=-17/26

OpenStudy (anonymous):

28x - 4 = 6x + 3 -4x - 24 28x - 4 = 2x - 21 28x -2x = -21 + 4 26x = -17 x = -17/26